# Diffusion in the Inverted Triangular Soft Lorentz Gas

### v1

@@ I Introduction
In this work, we investigate the inverted soft Lorentz gas, in which the repulsive potential bumps are replaced by attractive potential wells – similar to the experimental setup described in Ref. [20]. This modification leads to markedly different system behavior compared to the conventional square or triangular lattices with repulsive scatterers. In particular, our inverted setup exhibits an infinite horizon for all parameter regimes where particles possess enough energy to escape the wells and travel to infinity [21, 22, 23]. In soft-potential systems, variations in the potential landscape cause speed fluctuations along trajectories, leading to complex dynamics and regular structures in phase space [24, 25, 26]. An infinite horizon complicates analysis further; in hard-wall systems, it rules out normal diffusion [9, 10, 11], and similarly challenges conventional diffusion approximations in soft systems. In particular, the Machta–Zwanzig (MZ) approximation, which assumes uncorrelated hops between unit cells [21], breaks down in infinite-horizon lattices where normal diffusion is ill-defined.
@@ II Model and methodology
$V(\boldsymbol{r})=1+\sum_{n}V_{\mathrm{s}}(\boldsymbol{r}-\boldsymbol{r}_{n}),$ (2)
where $\boldsymbol{r}_{n}$ is the location of the center of the $n$th potential well and the ”+1” ensures that the potential is above zero in parameter space that we are interested in. For simplicity, we choose $V_{0}=1$. In Hartree atomic units ($r_{0}=m=1$), we set the total energy of the particle to $E=1/2$. As the potential height is at most one, we expect this energy choice to be somewhat representative of the dynamics in parameter space where the particles can cross between potential wells but cannot move entirely freely in the space. Furthermore, we define the unit cell side length as $L$, given by $L=2r_{0}+w$, where $w$ is the gap width between potential wells.
@@
with the angle brackets denoting a configurational ensemble average, forming the mean squared displacement (MSD) as $\langle(\boldsymbol{r}(t)-\boldsymbol{r}(0))^{2}\rangle$. Here, $\boldsymbol{r}(t)$ refers to the position of a particle at time $t$. We note that Eq. (3) is only valid for normal diffusion where the MSD grows linearly in time. Anomalous diffusion – superdiffusion and subdiffusion – is obtained when MSD $\propto t^{\alpha}$ with $\alpha>1$ and $\alpha<1$, respectively.
For computing the $D$ with Eq. (3), we require a sufficiently large ensemble of simulations. For this purpose, we employ the Bill2D software package [28]. We utilize the sixth-order symplectic integrator [29] with a timestep of $\Delta t=10^{-3}$, as this combination has been shown to produce the most accurate conservation of energy within the parameter space in our simulations. [15] To accurately compute the potential values affecting a particle, we form a 3-by-3 unit cell grid and place the particle in the centermost unit cell. Unless otherwise specified, we randomize the initial positions and propagation angles of the particles uniformly according to the energetically allowed phase space determined by the constant energy $E=1/2$.
@@ III.1 Parameter space
Due to the inverted nature of the potential, which is now characterized by attractive wells rather than repulsive bumps, the relevant regions of parameter space differ notably from those encountered in conventional Lorentz gas setups. This distinction is illustrated in Fig. 1. We can deduce that as a function of the gap width $w$ and system softness $\sigma$, several regions of differing behavior appear. First, when the potentials are far away from each other (large $w$) and/or the potential walls are hard (small $\sigma$), particles become confined within individual wells. In this limit, the system effectively behaves as a set of isolated soft circular billiards, with no diffusion. While we are not aware of direct studies of this specific configuration, Ref. [30] notes that soft elliptical billiards with low eccentricity lack a stochastic layer, which suggests the absence of chaotic dynamics. However, while not utilizing the exact Fermi potential, Ref. [31] suggests that even billiards with soft walls might exhibit chaos.
@@
This limit is denoted with the upper line in Fig. 2. This Figure also includes two dashed lines and one dotted line, representing areas of further analysis discussed in Sec. IVA.
@@ III.2 Approximations for $D$
In the following, we adopt the hopping model introduced in Ref. [15] to derive a more accurate expression for the diffusion coefficient. In this approach, trap-hopping sequences of length $n$ are considered, and the diffusion coefficient $D$ is computed based on the lengths and durations of these paths, weighted by their respective probabilities. Each hop is defined as an exit from a trap in a specific absolute direction. Our trap has six exits, and we define $\mathcal{H}=\{\nwarrow,\nearrow,\rightarrow,\searrow,\swarrow,\leftarrow\}$ as the alphabet of possible hopping directions. The direction of the $i$th hop is denoted with $\Delta_{i}\in\mathcal{H}$, and we can determine the displacement vector of a path $\Delta_{1}\Delta_{2}\ldots\Delta_{n}$ as
@@ IV.1 Diffusion coefficient
[…]
In Fig. 4, we also present $D_{\mathrm{MZ}}$ as a dash-dotted line, and note that it underestimates the diffusion coefficient even though the onset of diffusion is captured. Moreover, we present $D_{\mathrm{MZ}}^{n\mathrm{-hop}}$ for several values of $n$ as colored lines. Increasing the number of hops generally increases the value of the approximate diffusion coefficient. However, the approximation overshoots the actual diffusion coefficient, at least partly caused by the numerous COs in the lattice. When applying the correction in Eq. (17), presented as colored dashed lines, we see that the approximation becomes closer to the actual value of $D$, especially between $w=0.17-0.22$. We can also apply similar corrections to the computed diffusion coefficient by removing COs from the calculations. This result is shown as a dotted line in Fig. 4. Now, the corrected diffusion coefficient is close to the approximated value, especially between $w=0.17-0.22$.
@@ IV.2 Parameter space
[…]
Figure 5: (a) Density of quasiballistic trajectories ($\rho_{\mathrm{B}}$) in the parameter space of the inverted triangular soft Lorentz gas, characterized by a tongue-like structure inside a lagoon without quasiballistic trajectories. (b) Density of localized periodic orbits ($\rho_{\mathrm{LPO}}$) over the parameter space. LPOs are ubiquitous over almost the whole parameter space. The diffusion and free motion limits correspond to the lines in Fig. 2.
@@ IV.3 Example trajectories
In Fig. 7, we present some examples of LPOs present in the system. We observe a diverse range of LPOs, from trajectories confined to a single potential well, as illustrated in Fig. 7(f–m), to those spanning a few neighboring wells, shown in Fig. 7(a–e) and (o). Additionally, we identify highly intricate orbits that traverse multiple wells across the lattice, exemplified in Fig. 7(n) and (p). We simulated these trajectories to a maximum time of 40 0004000040\,00040 000, confirming the robustness of our algorithms and the stability of the trajectories.

### v2

@@ I Introduction
In this work, we investigate the inverted soft Lorentz gas, in which the repulsive potential bumps are replaced by attractive potential wells – similar to the experimental setup described in Ref. [20]. Moreover, triangular optical lattices with tunable potential landscapes have already been realized with ultracold atoms [21], further underlining the potential experimental relevance of our model. This modification leads to markedly different system behavior compared to the conventional square or triangular lattices with repulsive scatterers. In particular, our inverted setup exhibits an infinite horizon for all parameter regimes where particles possess enough energy to escape the wells and travel to infinity [22, 23, 24]. In soft-potential systems, variations in the potential landscape cause speed fluctuations along trajectories, leading to complex dynamics and regular structures in phase space [25, 26, 27]. An infinite horizon complicates analysis further; in hard-wall two-dimensional systems, it rules out normal diffusion [9, 10, 11], while in higher dimensions the situation is more complicated [24]. In soft systems such horizons similarly challenge conventional diffusion approximations. In particular, the Machta–Zwanzig (MZ) approximation, which assumes uncorrelated hops between unit cells [22], breaks down in infinite-horizon lattices where normal diffusion is ill-defined [15]. The main assumption behind the MZ approximation is that hops between traps are uncorrelated. In infinite horizon lattices, however, channels appear leading to particle movement along straight lines across (infinitely) many traps. This prevents randomisation in each trap, which leads to non-Markovian correlations in the dynamics.
@@ II Model and methodology
$V(\boldsymbol{r})={{V_{0}}}+\sum_{n}V_{\mathrm{s}}(\boldsymbol{r}-\boldsymbol{r}_{n}),$ (2)
where $\boldsymbol{r}_{n}$ is the location of the center of the $n$th potential well and the $V_{0}$” ensures that the potential is above zero in parameter space that we are interested in. For simplicity, we choose $V_{0}=1$. In Hartree atomic units ($r_{0}=m=1$), we set the total energy of the particle to $E=1/2$. As the potential height is at most one, we expect this energy choice to be representative of the dynamics in parameter space where the particles can cross between potential wells but cannot move entirely freely in the space. In the present work, we therefore varied only the gap width $w$ and the softness parameter $\sigma$ to access different dynamical regimes. We note, however, that varying the particle energy can also modify these regimes; see Ref. [29] for a detailed study. Furthermore, we define the unit cell side length as $L$, given by $L=2r_{0}+w$, where $w$ is the gap width between potential wells.
@@
with the angle brackets denoting a configurational ensemble average, forming the mean squared displacement (MSD) as $\langle(\boldsymbol{r}(t)-\boldsymbol{r}(0))^{2}\rangle$. Here, $\boldsymbol{r}(t)$ refers to the position of a particle at time $t$. We note that Eq. (3) is only valid for normal diffusion where the MSD grows linearly in time. Anomalous diffusion – superdiffusion and subdiffusion – is obtained when MSD $\propto t^{\alpha}$ with $\alpha>1$ and $\alpha<1$, respectively. The main sources of anomalous diffusion are COs, LPOs and quasiballistic orbits. We define quasiballistic orbits as regular trajectories that propagate infinitely toward one direction. However, we note that they may display periodic small-scale features along the way (see Fig. 8 below).
For computing the $D$ with Eq. (3), we require a sufficiently large ensemble of simulations. For this purpose, we employ the Bill2D software package [30]. We utilize the sixth-order symplectic integrator [31] with a timestep of $\Delta t=10^{-3}$, as this combination has been shown to produce the most accurate conservation of energy within the parameter space in our simulations. [15] To accurately compute the potential values affecting a particle, we form a 3-by-3 unit cell grid and place the particle in the centermost unit cell. Unless otherwise specified, as the invariant density of the system is not known, we randomize the initial positions and propagation angles of the particles uniformly according to the energetically allowed phase space determined by the constant energy $E=1/2$.
@@ III.1 Parameter space
Due to the inverted nature of the potential, which is now characterized by attractive wells rather than repulsive bumps, the relevant regions of parameter space differ notably from those encountered in conventional Lorentz gas setups. This distinction is illustrated in Fig. 1. We can deduce that as a function of the gap width $w$ and system softness $\sigma$, several regions of differing behavior appear. First, when the potentials are far away from each other (large $w$) and/or the potential walls are hard (small $\sigma$), particles become confined within individual wells. In this limit, the system effectively behaves as a set of isolated soft circular billiards slightly perturbed by the surrounding wells, and exhibits no diffusion. While we are not aware of direct studies of this specific configuration, Ref. [32] notes that soft elliptical billiards with low eccentricity lack a stochastic layer, which suggests the absence of chaotic dynamics. However, while not utilizing the exact Fermi potential, Ref. [33] suggests that even billiards with soft walls might exhibit chaos.
@@
This limit is denoted with the upper line in Fig. 2. This Figure also includes two dashed lines and one dotted line, representing areas of further analysis discussed in Sec. IVA. These areas were selected as representative slices through various parameter space structures further discussed in Sec. IVB.
@@ III.2 Approximations for $D$
In the following, we adopt the hopping model introduced in Ref. [15] to derive a more accurate expression for the diffusion coefficient. In this approach, trap-hopping sequences of length $n$ are considered, and the diffusion coefficient $D$ is computed based on the lengths and durations of these paths, weighted by their respective probabilities. Each hop is defined as an exit from a trap in a specific absolute direction. Our trap has six exits, and we define ${,,,,,}\mathcal{H}=\{\nwarrow$$\nearrow,\rightarrow$$\searrow,\swarrow$$\leftarrow\}$ as the alphabet of possible hopping directions. The direction of the $i$th hop is denoted with $\Delta_{i}\in\mathcal{H}$, and we can determine the displacement vector of a path $\Delta_{1}\Delta_{2}\ldots\Delta_{n}$ as
@@ IV.1 Diffusion coefficient
In Fig. 3(b) we present the diffusion coefficient as a function of system softness $\sigma$ at $w=0.1013$. In these simulations, N=40  [^0] N=\$40\,000\$ and T=40  [^0] T=\$40\,000\$ for each value of $\sigma$. We note that the diffusion coefficient has a complex dependency on the system parameters. Overall, it tends to increase with growing softness, as this leads to wider and shallower gaps between potential wells, facilitating particle transport. Notably, there are broad regions in parameter space where the diffusion coefficient is undefined due to the emergence of anomalous diffusion, caused by quasiballistic trajectories that violate the assumptions of normal diffusive behavior. These areas are marked with grey in the Figure.
In Fig. 4, the diffusion coefficient is presented as a function of $w$ for $\sigma=0.0989$. Once again, $D$ is a complex function of $w$. In the simulations, N=T=40  [^0] N=T=\$40\,000\$. Overall, as the spacing between wells increases, the diffusion coefficient tends to decline. This is caused by the narrowing gap where the particles can escape the well. Just before $w=0.22$, the gap closes, preventing any further diffusion. Furthermore, several gray regions indicate parameter ranges where $D$ is undefined due to the presence of quasiballistic trajectories that give rise to anomalous diffusion.
In Fig. 4, we also present $D_{\mathrm{MZ}}$ as a dash-dotted line, and note that it underestimates the diffusion coefficient even though the onset of diffusion is captured. Moreover, we present $D_{\mathrm{MZ}}^{n\mathrm{-hop}}$ for several values of $n$ as colored lines. Increasing the number of hops generally increases the value of the approximate diffusion coefficient. However, the approximation overshoots the actual diffusion coefficient, at least partly caused by the numerous COs in the lattice. When applying the correction in Eq. (17), presented as colored dashed lines, we see that the approximation becomes closer to the actual value of $D$, especially between w=0.170.22w=0.17-0.22. We can also apply similar corrections to the computed diffusion coefficient by removing COs from the calculations. This result is shown as a dotted line in Fig. 4. Now, the corrected diffusion coefficient is close to the approximated value, especially between w=0.170.22w=0.17-0.22.
However, despite correcting for COs, a slight disagreement between the numerical diffusion coefficient and the approximation remains. This is likely related to the complex phase-space structure with localized orbits, which hinders precise convergence. Thus, the approximation captures only the qualitative behavior, and the exact cause of the discrepancy is left for future work.
@@ IV.2 Parameter space
First, we study the proportion of quasiballistic trajectories ($\rho_{\mathrm{B}}$) in Fig. 5(a). In these simulations, N=100  [^0] N=\$100\,000\$ and T=1000T=\$1000\$. The dashed and dotted lines represent the free movement and diffusion thresholds, respectively, cf. Fig. 2. We did not conduct simulations in the white region at the top of the figure, as particles can move freely there, making the dynamics trivial from a diffusive perspective. Similarly, simulations were omitted below the threshold where particles are unable to escape the wells, since diffusion is entirely suppressed in that regime.
Figure 5: (a) Density of quasiballistic trajectories ($\rho_{\mathrm{B}}$) in the parameter space of the inverted triangular soft Lorentz gas, characterized by a tongue-like structure inside a lagoon without quasiballistic trajectories. The dashed vertical and horizontal lines refer to the parameter areas studied in Figs. 3(b) and 4, respectively. (b) Density of localized periodic orbits ($\rho_{\mathrm{LPO}}$) over the parameter space. LPOs are ubiquitous over almost the whole parameter space. The diffusion and free motion limits correspond to the lines in Fig. 2.
@@ IV.3 Example trajectories
In Fig. 7, we present some examples of LPOs present in the system. We observe a diverse range of LPOs, from trajectories confined to a single potential well, as illustrated in Fig. 7(f–m), to those spanning a few neighboring wells, shown in Fig. 7(a–e) and (o). Additionally, we identify highly intricate orbits that traverse multiple wells across the lattice, exemplified in Fig. 7(n) and (p). We simulated these trajectories to a maximum time of 40 00040\,000, confirming the robustness of our algorithms and the stability of the trajectories.